# Inflationary Dynamics and Perturbations in Fractal Cosmology

### v1

@@ 4 Perturbation Theory and Connecting with Observations
$f^{\prime\prime}+(-\nabla^{2}_{r}-\frac{z^{\prime\prime}}{z})f=0,$ (40)
@@
The Laplacian then transforms as
@@
$\nabla_{r}^{2}f=\left[a\frac{D}{3}\left(\frac{D}{3}-1\right)R^{1-\frac{6}{D}}(4\pi)^{2(D-3)/3D}L^{2(\frac{3}{D}-1)}k+\frac{D}{3}R^{1-\frac{3}{D}}L^{\frac{3}{D}-1}(4\pi)^{\frac{D-3}{3D}}k^{2}\right]f$ (46)
$=\left\{\frac{D}{3}k^{2}a^{2-\frac{6}{D}}(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}\left[\left(\frac{D}{3}-1\right)(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}k^{2(\frac{3}{D}-1)}+k^{\frac{3}{D}-1}a^{\frac{3}{D}-1}\right]\right\}f.$ (47)
@@
$f^{\prime\prime}+\left\{\frac{D}{3}k^{2}a^{2-\frac{6}{D}}(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}\left[\left(\frac{D}{3}-1\right)(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}k^{2(\frac{3}{D}-1)}+k^{\frac{3}{D}-1}a^{\frac{3}{D}-1}\right]-\frac{z^{\prime\prime}}{z}\right\}f=0.$ (48)
@@
$\omega^{2}=\frac{D}{3}k^{2}a^{2-\frac{6}{D}}(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}\left[\left(\frac{D}{3}-1\right)(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}k^{2(\frac{3}{D}-1)}+k^{\frac{3}{D}-1}a^{\frac{3}{D}-1}\right]-\frac{z^{\prime\prime}}{z}.$ (49)
@@
$k_{\text{eff}}^{2}=\frac{D}{3}k^{2}a^{2-\frac{6}{D}}(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}\left[\left(\frac{D}{3}-1\right)(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}k^{2(\frac{3}{D}-1)}+k^{\frac{3}{D}-1}a^{\frac{3}{D}-1}\right].$ (50)
@@
$D>3\sqrt{1-\frac{2}{\beta}}.$ (58)
@@
$D<3\sqrt{1-\frac{2}{\beta}}.$ (59)
But, in this case, $k_{\text{eff}}^{2}$ would undoubtedly be less than 0. This would mean that the evolution of $f$ during inflation is exponentially suppressed (the enhanced case is contradictory to the very nature of inflation) at sub-horizon scales, rather than oscillatory. All $D$ such that $D<3(1-\beta)$ will result in the sub-horizon evolution of $f$ being exponential instead of oscillatory. This includes $D$ such that $3\sqrt{1-\frac{2}{\beta}}<D<3(1-\beta)$ and hence, even for a subset of the solution of $D$ in the case of $2\epsilon_{1}+\epsilon_{2}>0$, $f$ will have an exponential behavior while for the case of $2\epsilon_{1}+\epsilon_{2}>0$, all solutions of $D$ will result in $f$ behave exponentially).
Now, since $D>3(1-\beta)>0$, $\beta$ must be less than 0. Moreover, for realistic values of the fractional length scale $L$ and the Hubble parameter $H$ during inflation, we see that $D$ must be less than $3$ and in particular, should satisfy $2.7\lesssim D\lesssim 3$, depending on the cosmological and inflationary model assumed.
Table 1: Plack 2018 Data Analysis
@@
Starobinsky ($R+R^{2})$ $D=2.7$ Slight tension with Planck Observation of $n_{s}$
Starobinsky ($R+R^{2})$ $D=2.8$ Agreement with Planck Observation of $n_{s}$
Starobinsky ($R+R^{2})$ $D=2.9$ Agreement with Planck Observation of $n_{s}$
Starobinsky ($R+R^{2})$ $D=3.0$ Agreement with Planck Observation of $n_{s}$
@@
A close examination of Table 1 reveals that, while the Starobinsky potential remains stable under moderate deviations from $D=3$, the most substantive phenomenological effect of the fractal framework appears in the Natural Inflation scenario. In conventional four-dimensional cosmology, compatibility with CMB measurements requires a super-Planckian decay constant, $f\gtrsim 5M_{Pl}$, which has long been regarded as a major theoretical obstacle for embedding Natural Inflation in a controlled ultraviolet completion [93, 119, 38, 129, 83]. In the fractal setting, however, the effective rescaling of the gravitational sector significantly alters this conclusion. The dependence of the slow-roll parameters on the factor $\left(\frac{2GL^{2}}{3\rho}\right)^{1-\frac{D}{3}}$ has a pronounced effect on Natural Inflation. For example, at $D=2.5$, the model admits values of the decay constant smaller than the Planck mass, a regime that is inaccessible in standard four-dimensional cosmology. This result may mitigate the traditional UV tension and hints towards the possibility that Natural Inflation can be reconciled with observational data within a parameter space that is inaccessible in the standard $D=3$ framework.
@@ 5 Conclusion
In this work, we have developed a systematic analysis of inflationary dynamics [59, 79, 3, 81] and perturbation theory [107, 12, 127] in the framework of fractal cosmology , where the effective dimensionality of spacetime is characterized by a non-integer parameter ($D$), emerging from a deformation of horizon thermodynamics and fractal measures [62, 27, 8, 104, 10, 89, 84, 91, 7, 106, 92, 29, 90, 76, 44, 13, 121, 22, 105]. The resulting Friedmann and Klein-Gordon equations acquire explicit $D$ dependent corrections, which in turn modify the slow-roll dynamics and the evolution of scalar perturbations.
Our work spanned three classes of inflationary models: a cubic potential, the Starobinsky ($R+R^{2}$) potential [93, 119, 38, 129, 83] and Natural inflationary potential [51, 50, 39]. In each case, the fractal dimension $D$ enters directly into the slow-roll parameters, the number of e-folds, and the form of the Mukhanov–Sasaki equation [107, 12, 127]. A distinctive feature of the fractal setup is the geometry induced suppression of slow-roll parameters, which acts in addition to the suppression due to potential flatness. As a result, inflation can be achieved even in models without a highly flattened potential, provided the effective dimension $D$ lies in an appropriate range.
A key phenomenological finding is that this geometric suppression reduces the unique observational advantage that plateau models like Starobinsky inflation [93, 119, 38, 129, 83] enjoy in standard four-dimensional cosmology. In the conventional scenario, the plateau potential of the Starobinsky model produces a strong flattening of the slow-roll parameters and an excellent fit to cosmic microwave background (CMB) data [19, 61, 35].In the fractal context, however, an additional suppression mechanism tied to the effective dimension partially replicates this effect, diminishing the importance of potential flatness alone. Consequently, Starobinsky inflation remains consistent with current CMB and baryon acoustic oscillation (BAO) constraints [9] for values $2.7\lesssim D\lesssim 3$, but it no longer stands out as uniquely preferred once fractal effects are included.
@@
Furthermore, Natural Inflation [51, 50, 39] in fractal cosmology exhibits significant phenomenological implications. The fractal corrections induce an effective rescaling of the Planck mass, which has the consequence of relaxing the conventional lower bound on the axion decay constant $f$. This may mitigate the well-known tension associated with the requirement of super-Planckian values of $f$ in standard four-dimensional Natural Inflation [93, 119, 38, 129, 83]. Within the fractal framework, Planck-compatible values of the spectral index allow substantially smaller decay constants, suggesting that Natural Inflation becomes easier to reconcile with ultraviolet completions.
At the perturbative level, we derived a fractal extension of the Mukhanov-Sasaki equation [107, 12, 127] by introducing an effective momentum $k_{\mathrm{eff}}$ which encodes the modification of the spatial Laplacian induced by the fractal geometry. This leads to explicit $D$ and $L$ dependent corrections to the scalar power spectrum and to the spectral index $n_{s}$. By comparing our predictions with the Planck 2018 constraint on the scalar spectral index, $n_{s}=0.9649\pm 0.0042$ [9], we find that the effective fractal dimension is tightly bounded around the standard value ($D=3$), with a model-dependent allowed range ($2.7\lesssim D\lesssim 3$). In the natural inflationary case, decreasing the value of $D$ effectively decreases the lower bound of the axion decay constant $f$.
@@
Overall, our analysis demonstrates that fractal cosmology [62, 27, 8, 104, 10, 89, 84, 91, 7, 106, 92, 29, 90, 76, 44, 13, 121, 22, 105] offers a novel extension of the inflationary paradigm [59, 79, 3, 81]. There are several natural extensions of this work. One immediate direction is to study models like warm inflation [20, 21, 68, 15] in the fractal setup, phases like ultra slow roll [41, 86] and modified background cosmologies like RSII braneworld [82, 75, 24, 114]. Another possible direction is to study reheating [14, 70, 71, 4] (instead of avoiding it by considering warm inflation) in the fractal framework and to determine how the fractal dimension modifies the efficiency of energy transfer at the end of inflation.
@@ 6 Acknowledgments
The work of M. K. was performed in Southern Federal University with financial support of grant of Russian Science Foundation № 25-07-IF. P.M. acknowledges the FCT grant UID/212/2025 Centro de Matem´atica e Aplica¸c˜oes da Universidade da Beira Interior plus the COST Actions CA23130 (Bridging high and low energies in search of quantum gravity (BridgeQG)) and CA23115 (Relativistic Quantum Information (RQI)). The work of O.T. was supported in part by the Vanderbilt Discovery Doctoral Fellowship. We are grateful to Prof. Pankaj Joshi, Meet Vyas, Aum Trivedi and peers at ICSC for their support.

### v2

@@ 4 Perturbation Theory and Connecting with Observations
$f^{\prime\prime}+\left(-\nabla^{2}_{r}-\frac{z^{\prime\prime}}{z}\right)f=0,$ (41)
@@
The Laplacian  [^4] There exist other works in literature who have, in a similar spirit, studied the effect of fractional modifications to the Laplacian and d’Alembertian. The work by El-Nabulsi & Anukool [40] introduces functionality through Dunkl fractional Laplacian operators. Furthermore, Calcagni and Rachwał have explored UV complete quantum field theories with a fractional d’Alembertian [19] then transforms as
@@
$\begin{split}\nabla_{r}^{2}f=\frac{D}{3}k^{2}a^{2-\frac{6}{D}}(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}\\ \left[\left(\frac{D}{3}-1\right)(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}k^{2(\frac{3}{D}-1)}+k^{\frac{3}{D}-1}a^{\frac{3}{D}-1}\right]f.\end{split}$ (47)
@@
$\begin{split}f^{\prime\prime}+\\ \left\{\begin{split}\frac{D}{3}k^{2}a^{2-\frac{6}{D}}(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}\left[\left(\frac{D}{3}-1\right)(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}k^{2(\frac{3}{D}-1)}+k^{\frac{3}{D}-1}a^{\frac{3}{D}-1}\right]\\ -\frac{z^{\prime\prime}}{z}\end{split}\right\}f=0.\end{split}$ (48)
@@
$\begin{split}\omega^{2}=\frac{D}{3}k^{2}a^{2-\frac{6}{D}}(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}\\ \left[\left(\frac{D}{3}-1\right)(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}k^{2(\frac{3}{D}-1)}+k^{\frac{3}{D}-1}a^{\frac{3}{D}-1}\right]-\frac{z^{\prime\prime}}{z}.\end{split}$ (49)
@@
$\begin{split}k_{\text{eff}}^{2}=\frac{D}{3}k^{2}a^{2-\frac{6}{D}}(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}\\ \left[\left(\frac{D}{3}-1\right)(4\pi)^{\frac{D-3}{3D}}L^{\frac{3}{D}-1}k^{2(\frac{3}{D}-1)}+k^{\frac{3}{D}-1}a^{\frac{3}{D}-1}\right].\end{split}$ (50)
@@
$D>3\left(\frac{2-\beta}{\beta+2}\right).$ (58)
@@
$D<3\left(\frac{2-\beta}{\beta+2}\right).$ (59)
In this regime the effective momentum satisfies, $k_{\text{eff}}^{2}<0$ implying that $f$ evolves exponentially on sub-horizon scales rather than exhibiting the usual oscillatory behavior. Such exponential evolution is incompatible with the standard inflationary picture, where sub-horizon modes must oscillate in order to define a consistent quantum vacuum. Consequently, all values of the effective dimension satisfying
$D<3(1-\beta)$ (60)
lead to an unphysical sub-horizon dynamics of scalar perturbations. Requiring the dynamics to remain physical therefore imposes $D>3(1-\beta)>0$, which immediately implies $\beta<1$. For phenomenologically reasonable values of the fractional length scale $L$ and the Hubble parameter $H$ during Starobinsky ($R+R^{2}$) inflation, this condition restricts the effective fractal dimension to lie close to the standard three-dimensional value. In practice, consistency with oscillatory sub-horizon modes and observational considerations suggests a best-fit range of $2.7\lesssim D\lesssim 3$ for Starobinsky ($R+R^{2}$) inflation .
Table 1: Planck 2018 Data Analysis
@@
Starobinsky ($R+R^{2})$ $D=2.7$ $n_{s}\simeq 0.961616$
Starobinsky ($R+R^{2})$ $D=2.8$ $n_{s}\simeq 0.962338$
Starobinsky ($R+R^{2})$ $D=2.9$ $n_{s}\simeq 0.963009$
Starobinsky ($R+R^{2})$ $D=3.0$ $n_{s}\simeq 0.963636$
@@
A close examination of Table 1 reveals that, while the Starobinsky potential remains stable under moderate deviations from $D=3$, the most substantive phenomenological effect of the fractal framework appears in the Natural Inflation scenario. In conventional four-dimensional cosmology, compatibility with CMB measurements requires a super-Planckian decay constant, $f\gtrsim 5M_{Pl}$, which has long been regarded as a major theoretical obstacle for embedding Natural Inflation in a controlled ultraviolet completion [48, 47, 35]. In the fractal setting, however, the effective rescaling of the gravitational sector significantly alters this conclusion. The dependence of the slow-roll parameters on the factor $\left(\frac{2GL^{2}}{3\rho}\right)^{1-\frac{D}{3}}$ has a pronounced effect on Natural Inflation. For example, at $D=2.5$, the model admits values of the decay constant smaller than the Planck mass, a regime that is inaccessible in standard four-dimensional cosmology. This result may mitigate the traditional UV tension and hints towards the possibility that Natural Inflation can be reconciled with observational data within a parameter space that is inaccessible in the standard $D=3$ framework.
More broadly, this illustrates how the fractal framework can modify the phenomenological viability of inflationary models rather than simply shifting numerical predictions. In particular, the modification of the slow-roll relations changes the mapping between the parameters of the potential and observable quantities such as the scalar spectral index. As a consequence, regions of parameter space that are excluded in conventional cosmology can become compatible with observational constraints. The Natural Inflation scenario provides a clear example of this effect: the fractal modification effectively enlarges the viable parameter space by relaxing the lower bound on the decay constant $f$. In this sense, the fractal framework provides a concrete phenomenological advantage, as it allows models that are otherwise theoretically disfavored due to super-Planckian field excursions to remain compatible with current cosmological observations.
@@ 5 Conclusion
In this work, we have developed a systematic analysis of inflationary dynamics [53, 66, 3, 68] and perturbation theory [90, 11, 106] in the framework of fractal cosmology , where the effective dimensionality of spacetime is characterized by a non-integer parameter ($D$), emerging from a deformation of horizon thermodynamics and fractal measures [55, 21, 7, 86, 9, 74, 69, 76, 6, 89, 77, 25, 75, 64, 38, 12, 101, 15, 16, 88, 32, 52]. The resulting Friedmann and Klein-Gordon equations acquire explicit $D$ dependent corrections, which in turn modify the slow-roll dynamics and the evolution of scalar perturbations.
Our work spanned four classes of inflationary models: a linear potential, a cubic potential, the Starobinsky ($R+R^{2}$) potential [99] and Natural inflationary potential [48, 47, 35]. In each case, the fractal dimension $D$ enters directly into the slow-roll parameters, the number of e-folds, and the form of the Mukhanov-Sasaki equation [90, 11, 106]. A distinctive feature of the fractal setup is the geometry induced suppression of slow-roll parameters, which acts in addition to the suppression due to potential flatness. As a result, inflation can be achieved even in models without a highly flattened potential, provided the effective dimension $D$ lies in an appropriate range.
A key phenomenological finding is that this geometric suppression reduces the unique observational advantage that plateau models like Starobinsky inflation [99] enjoy in standard four-dimensional cosmology. In the conventional scenario, the plateau potential of the Starobinsky model produces a strong flattening of the slow-roll parameters and an excellent fit to cosmic microwave background (CMB) data [14, 54, 31].In the fractal context, however, an additional suppression mechanism tied to the effective dimension partially replicates this effect, diminishing the importance of potential flatness alone. Consequently, Starobinsky inflation remains consistent with the Planck 2018 observations for values $2.7\lesssim D\lesssim 3$, but it no longer stands out as uniquely preferred once fractal effects are included.
@@
Furthermore, Natural Inflation [48, 47, 35] in fractal cosmology exhibits significant phenomenological implications. The fractal corrections induce an effective rescaling of the Planck mass, which has the consequence of relaxing the conventional lower bound on the axion decay constant $f$. This may mitigate the well-known tension associated with the requirement of super-Planckian values of $f$ in standard four-dimensional Natural Inflation [48, 47, 35]. Within the fractal framework, Planck-compatible values of the spectral index allow substantially smaller decay constants, suggesting that Natural Inflation becomes easier to reconcile with ultraviolet completions.
At the perturbative level, we derived a fractal extension of the Mukhanov-Sasaki equation [90, 11, 106] by introducing an effective momentum $k_{\mathrm{eff}}$ which encodes the modification of the spatial Laplacian induced by the fractal geometry. This leads to explicit $D$ and $L$ dependent corrections to the scalar power spectrum and to the spectral index $n_{s}$. By comparing our predictions with the Planck 2018 constraint on the scalar spectral index, $n_{s}=0.9649\pm 0.0042$ [8], we find that the effective fractal dimension has a best-fit range $2.7\lesssim D\lesssim 3$ for Starobinsky ($R+R^{2}$) inflation. In the natural inflationary case, decreasing the value of $D$ effectively decreases the lower bound of the axion decay constant $f$.
@@
Overall, our analysis demonstrates that fractal cosmology [55, 21, 7, 86, 9, 74, 69, 76, 6, 89, 77, 25, 75, 64, 38, 12, 101, 16, 15, 88, 52, 32] offers a novel extension of the inflationary paradigm [53, 66, 3, 68].
@@ 6 Acknowledgments
The work of M. K. was performed in Southern Federal University with financial support of grant of Russian Science Foundation № 25-07-IF. P.M. acknowledges the FCT grant UID/212/2025 CMA-UBI plus the COST Actions CA23130 (Bridging high and low energies in search of quantum gravity (BridgeQG)) and CA23115 (Relativistic Quantum Information (RQI)). The work of O.T. was supported in part by the Vanderbilt Discovery Doctoral Fellowship. We are grateful to Prof. Pankaj Joshi, Meet Vyas, Aum Trivedi and peers at ICSC for their support. We also thank Prof. Luigi and Prof. Gopintah Guin for helpful comments and insightful discussions.